OpenStudy (anonymous):

A rectangle of length 10 meters and width 4 meters is inscribed in a circle. How long is the radius of the circle?

OpenStudy (anonymous):

draw the picture u will get the answer easily

OpenStudy (imstuck):

It looks like this:

OpenStudy (imstuck):



OpenStudy (imstuck):

Your radius is what you need to find, and consequently, it is also half the length of the diagonal of the rectangle. Let's look at it a different way now:

OpenStudy (imstuck):

When we pull that rectangle out using the diagonal as a boundary, we have a right triangle with height of 4 and base of 10 and the hypotenuse is the diagonal of the rectangle. So we need to find the length of the hypotenuse and divide it in half and that's the length of the radius of the circle.

OpenStudy (imstuck):



OpenStudy (imstuck):

Using Pythagorean's theorem we know that x^2 will equal 4^2 + 10^2, which is

OpenStudy (imstuck):

\[x ^{2}=16+100\]and \[x=\sqrt{116} \]or x = 10.770

OpenStudy (imstuck):

Dividing that in half you will have half the diagonal which is also the radius. 10.770/2=5.385

OpenStudy (imstuck):

Ty for the medal!

OpenStudy (anonymous):

Thanks!

OpenStudy (imstuck):

You bet!

OpenStudy (anonymous):

Is it necessary to be in decimal?